From GlueXWiki
Jump to: navigation, search

April 30, 2015 FDC+CDC meeting


  1. Instructions for Bluejeans meeting connection
  2. FDC meeting ID: 290664653
  3. To join via a Web Browser, go to the page [1] https://bluejeans.com/290664653.


  1. Collaboration meeting talks (meeting page)
  2. Gas puzzle
    • CO2 monitoring [2]
  3. CDC update (Beni, Mike)
  4. FDC update (Lubomir)
  5. Electronics (Fernando, Cody, Chris, Nick)
  6. Engineering
  7. Other


Participants: Curtis (CMU), Sean (NU), Luke, Fernando, Mike, Dave, Chris, Nick, Beni, Simon, Eugene, and Lubomir (JLab). Short meeting during the last week of the spring 2015 run.

Collaboration meeting talks

- Everybody at the meeting agreed with the agenda for the collaboration meeting.

Gas puzzle

- Last CO2 results linked above. Last days, the sensor was connected to the first FDC package. It is strange that we see now about 59% CO2, while before it was ~55%. The only change we did intentionally is the change of the CO2 gas factor from 0.77 to 0.74, which would correspond to 55->56% change of the CO2.

CDC update

- Mike: CDC is working fine. One HV channel (G14) is not stable, most likely problem in the CAEN module.

FDC update

- First look at the FDC efficiency now with 6.3V threshold on the discriminators. Compared to the 8V thresholds (on the picture attached: black 6.3V, red 8V from previous run) before we see just a little increase of the efficiency from 94 to about 95%: not significant; could also be noise.


- The power supply of a NIM crate in EEL126 used in the make a cosmic trigger failed; people there smelled it burning. The crate was labeled. Nick and Chris will look into this, and possibly replace the power supply.


- Tests with the spare package in the tagger. The chambers see high rates from: 1)background photons over the whole sensitive area and 2) electrons from the radiator in the shear plane of the magnet of the order of 100kHz per wire. Electrons, unlike photons, make hits in the two chambers of the spare package, that's how they are separated.